# Joint multi-band deconvolution for Euclid and Vera C. Rubin images

### v1

@@ 2.5 Gradient descent step-size
$C_{\{r,i,z\}}\geq\frac{\mathbf{h}^{\top}_{\{r,i,z\}}\ast\mathbf{h}_{\{r,i,z\}}}{\left\|\mathbf{\sigma}_{\{r,i,z\}}\right\|_{F}^{2}}+2\lambda_{\{r,i,z\}}\alpha^{2}_{\{r,i,z\}}\mathbf{h}_{euc}^{\top}\ast\mathbf{h}_{euc}$ (15)
Once that is found, the optimal constraints on the individual step-sizes $\beta_{\{r,i,z\}}$ that ensure convergence are as follows:
$\beta_{\{r,i,z\}}$ $\leq\frac{1}{C_{\{r,i,z\}}}$ (16)
From equations 15 and 16, it is important to note that the step-size also depends on the $Rubin$ and $Euclid$ PSFs (detailed in Sections 3.2 and 3.3), the values of $\lambda_{\{r,i,z\}}$ (described in Section 2.3) and $\alpha_{\{r,i,z\}}$ (shown in Section 2.2).
@@ 3.1 Ground truth images
We extracted HST cutout windows of dimensions $128\times 128$ pixels from GOODS-N and GOODS-S (Retzlaff, J. et al., 2010) in the $F606W$, $F775W$, and $F850LP$ bands by centering them at the centroid of the object. These HST bands were selected because their central wavelengths align with those of the $Rubin$ $r$-, $i$-, and $z$-bands, and these HST images were subsequently used to simulate the Rubin images, as explained in Section 3.2. The mosaiced HST ACS images along with the catalogue can be found at this link [^1] https://archive.stsci.edu/prepds/goods/. These HST cutouts are at a pixel scale of $0.05\arcsec$. We aimed to perform the experiments on galaxies with sizes large enough to effectively assess the impact of deconvolution on our ability to resolve their structural and morphological features, such as arms, bars, and clumps. To ensure the selection of large galaxies and exclude point-sized objects, we applied the following filtering criteria to the $F775W$ band catalogue:
@@
We then visually inspected and selected 92 objects that exhibited extended and complex galaxy structures. The histogram of the HST $F775W$ band magnitude for all galaxies in our dataset is shown in Figure 2. It is important to highlight that according to our simulations, the HST $F775W$ band corresponds to the Rubin $i$-band, and we will refer to it as the Rubin $i$-band throughout the text.
Figure 2: Histogram of the HST $F775W$ band magnitude for all galaxies in our dataset after filtering. Note that the HST $F775W$ band matches with the Rubin $i$-band.
@@
The simulated Vera C. Rubin-like PSFs were obtained from the second data challenge (DC2) of the Legacy Survey of Space and Time (LSST) Dark Energy Science Collaboration (DESC). The atmospheric and optical effects, as well as sensor-induced electrostatic effects, are simulated using physically motivated models, with a final adjustment to the PSF sizes to match the expected data from the Rubin Observatory, as described in detail in Abolfathi et al. (2021).
To generate the Vera C. Rubin simulated images, we first convolved the HST images in the $F606W$, $F775W$, and $F850LP$ bands with the Rubin $r$-, $i$-, and $z$-band PSFs, respectively, such that the simulated images are at the expected Rubin resolution, with a pixel scale of $0.2\arcsec$. Subsequently, we added white Gaussian noise such that our Rubin-simulated images have a signal-to-noise (S/N) ratio ranging between $12$ and $28$, with a median around 20. Based on the survey parameters outlined in Željko Ivezić et al. (2019), our simulations suggest an S/N range that corresponds to the initial few visits of the telescope. This indicates that our method could be effectively applied as soon as the first images start arriving.
@@ 3.3 Euclid images
The simulated Euclid-like $VIS$-band PSF was obtained using the WaveDiff model proposed by Liaudat et al. (2022), which changes the data-driven PSF modelling space from the pixels to the wavefront by adding a differentiable optical forward model in the modelling framework. WaveDiff outputs an approximation of the true Euclid PSF, which was derived before the actual launch of the satellite.
Next, we calculated the fractional flux contributions $\alpha_{r},\alpha_{i},\alpha_{z}$ by integrating the area under the curves in Figure 1 and normalising them to sum to 1. The resulting values, as also mentioned in Section 2.2, are $\alpha_{r}=0.3785$, $\alpha_{i}=0.3468$, and $\alpha_{z}=0.2746$. To generate the Euclid simulated images, we multiplied $\alpha_{r},\alpha_{i},\alpha_{z}$ with the HST images in the $F606W$, $F775W$, and $F850LP$ bands, respectively. The result is then convolved with the simulated Euclid PSF, such that the simulated images are at the expected Euclid resolution with a pixel scale of $0.1\arcsec$. Finally, white Gaussian noise is added such that the Euclid-simulated images have a signal-to-noise (S/N) ratio ranging between $20$ and $45$, with a median around 35. Based on the calculations presented by Euclid Collaboration et al. (2022) which assess the S/N statistics for Euclid, our simulations are conservative, implying that our method would perform well when applied to real images with higher S/N.
@@ 4 Results
The algorithm simultaneously processed the noisy simulations from the three Vera C. Rubin bands and the Euclid $VIS$-band, along with their respective PSFs. These noisy images served as initializations or first guesses for the algorithm. Subsequently, the algorithm iteratively minimised the loss functions 6-8, as detailed in Section 2.4. The algorithm was run for 200 iterations, with convergence typically observed within $50$-$100$ iterations for all images in our dataset. Figure 3 shows the convergence plot of the loss function for the deconvolved output in Figure LABEL:subfig:mcdec1.
Figure 3: The loss function for the galaxy shown in Figure LABEL:subfig:mcdec1. Convergence is guaranteed at around 100 iterations when the relative change in loss value $<10^{-3}$ and the curve is flat.
@@ 4.1 Flux leakage test
As a validation to ensure no flux leakage between channels during the joint deconvolution of Vera C. Rubin and Euclid images, we conducted a unit test. We assumed three distinct Gaussians placed separately in the Rubin channels, with the Euclid simulation being a weighted sum of these three images, resulting in three disjoint Gaussians. Post-deconvolution, the Gaussians remained intact without any structure extending beyond their boundaries. This confirms that the structure present in the deconvolved image within each Rubin band is independent of the structures in other bands and each image accurately retains only the information relevant to its specific band. Figure 4 provides a visual demonstration of these findings.
@@ 4.2 Deconvolved outputs
We present two examples of deconvolved images in Figure LABEL:fig:deconv_outputs, illustrating the algorithm’s capability to recover features that were lost in the original Vera C. Rubin simulations. Visually, the deconvolved outputs seem to capture the variations between different bands. Qualitatively, these outputs exhibit high quality with minimal background noise and result in clean residuals. However, in the next Section, 4.3, we demonstrate that employing a deep learning-based denoiser further enhances the already impressive results achieved initially. Table 2 presents the normalized mean squared error (NMSE) computed with respect to the ground-truth HST image for the pre-denoised and post-deonised images.
4.3 Deep learning-based denoising
@@
After obtaining the deconvolved outputs, we feed them to DRUNet, a neural network proposed by Zhang et al. (2022) that combines U-Net (Ronneberger et al., 2015) and ResNet (He et al., 2016). U-Net is renowned for its efficiency in image-to-image translation, while ResNet excels in increasing modelling capacity through stacked residual blocks. Inspired by FFDNet (Zhang et al., 2018a), which incorporates a noise level map as input, DRUNet enhances U-Net by integrating residual blocks to improve prior denoising modelling. Similar approaches that combine U-Net and ResNet can be found in other studies (Zhang et al., 2018b; VenkateshG et al., 2018). The backbone of DRUNet is a U-Net architecture with four scales, and the schematic proposed by Zhang et al. (2022) is depicted in Figure 5.
@@
We chose DRUNet because it is a non-blind denoiser, meaning it takes the noise map as input. This ensures that any unknown noise can be estimated and given as input for denoising. We estimate the noise level map in our deconvolved images using the scikit-image package (van der Walt et al., 2014) by calculating the average noise level within four $16\times 16$ pixel squares placed at each corner of the image. This approach ensures that only background noise is measured, avoiding any contribution from the signal. This map is then fed to the pre-trained DRUNet, along with the deconvolved image. The denoised outputs are illustrated in Figure LABEL:fig:denoise_outputs. It is observed that the denoiser exclusively eliminates noise from the background without affecting the main structure of the galaxies. Although the enhancement in image quality is marginal (since the original deconvolved image is already of high quality), the normalized mean squared error (NMSE) computed with respect to the ground-truth HST image decreases, as shown in Table 2. The most notable improvement is observed in the $z$-band images. Even though DRUNet was trained on a combination of images from BSD (Chen & Pock, 2017), Waterloo Exploration Database (Ma et al., 2017), DIV2K (Agustsson & Timofte, 2017), and Flick2K (Lim et al., 2017), it is remarkable that it works well on astronomical data, showing great generalization. Finally, the fractional error in the output flux as a function of the $i$-band magnitude is shown in Figure 6, which indicates that the mean flux error is less than $5\%$ for the entire magnitude range for all the bands.
@@
Figure 5: The DRUNet architecture, which incorporates an additional noise level map as input and integrates U-Net (Ronneberger et al., 2015) with ResNet (He et al., 2016).”SConv” stands for strided convolution and ”TConv” stands for transposed convolution. Image credits: Zhang et al. (2022).
@@
Figure 6: Fractional error in the output flux as a function of the $i$-band magnitude (which is chosen in order to have the same scale on the x-axes). The dots correspond to the individual galaxies, and the grey line is the best-fit line after binning the magnitude values.
@@
NMSE $r$-band $i$-band $z$-band
@@
 [^2] NMSE is calculated for the pre-denoised and post-denoised images with respect to the ground-truth HST image.
4.4 Improvement compared to independent deconvolution of each band
To demonstrate the advantages of jointly deconvolving multiple photometric bands, we performed independent deconvolution of each band using the deconvolution framework introduced in Akhaury et al. (2024). The method proceeds in two steps: it first deconvolves the input using Tikhonov regularization (Tikhonov & Arsenin, 1977), and denoises the result using SUNet (Fan et al., 2022), a state-of-the-art Swin transformer-based architecture. The results, presented in Figure 8, show that the joint deconvolution method outperforms the independent deconvolution of individual bands. The joint method enables us to leverage the correlation between the different bands and the space-based image, thus improving the final output.
@@ 5 Conclusion
In this work, we have presented a novel image deconvolution technique designed to improve the resolution of multi-band ground-based data by leveraging higher-resolution space-based observations. Our approach, which focusses on the joint deconvolution of Vera C. Rubin and Euclid images, effectively exploits the overlapping spectral coverage of the Rubin $r$-, $i$-, and $z$-bands with the Euclid $VIS$-band. Through rigorous testing, we demonstrated that our iterative algorithm successfully recovers fine details while preserving the flux for each band. Different noise levels were tested, and the resolution achieved by ground-based data is close to that of HST. Our results indicate that joint deconvolution of all data outperforms independent deconvolution of individual photometric bands using existing state-of-the-art methods. By utilising the Euclid $VIS$-band image as a term that provides additional information, our technique ensures that the deconvolved Rubin images retain high spatial resolution and accurate photometric measurements. The integration of deep learning-based denoising using DRUNet enhances the quality of the deconvolved output, further reducing background noise without altering the main structures of the galaxies.
The potential applications of our method extend beyond the Euclid-Rubin pair, offering a versatile solution to improve the resolution of ground-based images in multiple photometric bands as long as there exists a space-based image of the same field of view in a band that encompasses all ground-based filters. In the future, we intend to test our deconvolution method on images of the Perseus cluster by using ground-based observations from the CFHT (Canada–France–Hawaii Telescope) and space-based observations from the Euclid ERO (Early Release Observations) public release.
6 Data availability
@@ Acknowledgements.
Figure 8: The galaxy in Figure LABEL:subfig:mcdec2 deconvolved using two different approaches, illustrating that joint deconvolution outperforms independent deconvolution of individual photometric bands. The joint method allows us to leverage the correlation between the different bands and the space-based image, thus improving the final output. The first column displays the Euclid $VIS$ image, followed by Vera C. Rubin simulations in the $r$, $i$, and $z$ bands in the second column. The independently deconvolved SUNet outputs for the three bands are shown in the third column. The fourth column shows the corresponding joint deconvolution outputs followed by denoising with DRUNet. The ground-truth HST images are shown in the fifth column.
Appendix A Generalization to objects with non-flat spectral energy distribution (SED)
To assess the impact of information transfer across different bands for objects with a non-flat spectral energy distribution (SED), we conducted an experiment where we successively replaced the Rubin bands with pure noise, ensuring that no galaxy signal was present. We then analyzed how features from the high-resolution Euclid image propagated into the deconvolved Rubin bands. The outputs are shown in Figure LABEL:fig:noise_test.
Our results confirm that reconstructed features appear only in bands where the original galaxy signal is present. This demonstrates that the algorithm does not artificially imprint Euclid information onto Rubin bands lacking real data. In cases where a structure is visible in the Euclid image but absent from one or more Rubin bands, the problem becomes degenerate: the feature could either be entirely attributed to a single band or distributed across multiple bands. However, our findings indicate that the outputs are directly influenced by the input data in each band rather than being dictated solely by the high-resolution Euclid image. Our results demonstrate that the joint deconvolution method effectively utilises the available signal in each band while respecting the constraints imposed by the data. This also confirms that the method would work for objects with non-flat SEDs, where signal may be present in only one band but absent in others, ensuring that features are accurately transferred according to their actual distribution across the bands.
As mentioned in Section 4, we found that the algorithm converges within $200$ iterations when the signal is available in all three bands. When one band is replaced with a noisemap, as shown in Figure LABEL:subfig:noise_test1, convergence requires approximately $1000$ iterations. Replacing a second band with a noisemap, as seen in Figure LABEL:subfig:noise_test2, further increases the iteration count to around $5000$. This demonstrates that incorporating more data across different bands significantly accelerates the convergence of the loss functions as the algorithm can benefit by capturing the correlations between these bands.

### v2

@@ 2.5 Gradient descent step size
$C_{\{r,i,z\}}\geq\frac{\mathbf{h}^{\top}_{\{r,i,z\}}\ast\mathbf{h}_{\{r,i,z\}}}{\left\|\mathbf{\sigma}_{\{r,i,z\}}\right\|_{F}^{2}}+2\lambda_{\{r,i,z\}}\alpha^{2}_{\{r,i,z\}}\mathbf{h}_{euc}^{\top}\ast\mathbf{h}_{euc}.$ (15)
Once that is found, the optimal constraints on the individual step sizes, $\beta_{\{r,i,z\}}$, that ensure convergence are as follows:
$\beta_{\{r,i,z\}}$ $\leq\frac{1}{C_{\{r,i,z\}.}}$ (16)
From Eqs. 15 and 16, it is important to note that the step size also depends on the Rubin and Euclid PSFs (see details in Sects. 3.2 and 3.3) and the values of $\lambda_{\{r,i,z\}}$ (described in Sect. 2.3) and $\alpha_{\{r,i,z\}}$ (shown in Sect. 2.2).
@@ 3.1 Ground truth images
We extracted HST cutout windows of dimensions $128\times 128$ pixels from GOODS-N and GOODS-S (Retzlaff, J. et al., 2010) in the $F606W$, $F775W$, and $F850LP$ bands by centering them at the centroid of the object. These HST bands were selected because their central wavelengths align with those of the Rubin $r$, $i$, and $z$ bands, and these HST images were subsequently used to simulate the Rubin images, as explained in Sect. 3.2. The mosaicked HST ACS (Advanced Camera for Surveys) images along with the catalog can be found at this link [^1] https://archive.stsci.edu/prepds/goods/. These HST cutouts are at a pixel scale of $0.05\arcsec$. We aimed to perform the experiments on galaxies with sizes large enough to effectively assess the impact of deconvolution on our ability to resolve their structural and morphological features, such as arms, bars, and clumps. To ensure the selection of large galaxies and exclude point-sized objects, we applied the following filtering criteria to the $F775W$ band catalog:
@@
We then visually inspected and selected 92 objects that exhibited extended and complex galaxy structures. The histogram of the HST $F775W$ band magnitude for all galaxies in our dataset is shown in Fig. 2. It is important to highlight that, according to our simulations, the HST $F775W$ band corresponds to the Rubin $i$ band, and we refer to it as the Rubin $i$ band throughout the text.
Figure 2: Histogram of the HST $F775W$-band magnitude for all galaxies in our dataset after filtering. Note that the HST $F775W$ band matches with the Rubin $i$ band.
@@
The simulated Rubin-like PSFs were obtained from the second data challenge (DC2) of the Legacy Survey of Space and Time (LSST) Dark Energy Science Collaboration (DESC). The atmospheric and optical effects, as well as sensor-induced electrostatic effects, are simulated using physically motivated models, with a final adjustment to the PSF sizes to match the expected data from the Rubin Observatory, as described in detail in Abolfathi et al. (2021).
To generate the Rubin simulated images, we first convolved the HST images in the $F606W$, $F775W$, and $F850LP$ bands with the Rubin $r$-, $i$-, and $z$-band PSFs, respectively, such that the simulated images are at the expected Rubin resolution, with a pixel scale of $0.2\arcsec$. Subsequently, we added white Gaussian noise such that our Rubin-simulated images have a S/N ranging between $12$ and $28$, with a median around 20. Based on the survey parameters outlined in Željko Ivezić et al. (2019), our simulations suggest an S/N range that corresponds to the initial few visits of the telescope. This indicates that our method could be effectively applied as soon as the first images start arriving.
@@ 3.3 Euclid images
The simulated Euclid-like VIS-band PSF was obtained using the WaveDiff model proposed by Liaudat et al. (2022), which changes the data-driven PSF modeling space from the pixels to the wavefront by adding a differentiable optical forward model in the modeling framework. WaveDiff outputs an approximation of the true Euclid PSF, which was derived before the actual launch of the satellite.
Next, we calculated the fractional flux contributions $\alpha_{r},\alpha_{i},\alpha_{z}$ by integrating the area under the curves in Fig. 1 and normalizing them to sum to 1. The resulting values, as also mentioned in Sect. 2.2, are $\alpha_{r}=0.3785$, $\alpha_{i}=0.3468$, and $\alpha_{z}=0.2746$. To generate the Euclid simulated images, we multiplied $\alpha_{r},\alpha_{i}$, and $\alpha_{z}$ by the HST images in the $F606W$, $F775W$, and $F850LP$ bands, respectively. The result is then convolved with the simulated Euclid PSF such that the simulated images are at the expected Euclid resolution with a pixel scale of $0.1\arcsec$. Finally, white Gaussian noise is added such that the Euclid-simulated images have a S/N ranging between $20$ and $45$, with a median around 35. Based on the calculations presented by Euclid Collaboration et al. (2022), who assessed the S/N statistics for Euclid, our simulations are conservative, implying that our method would perform well when applied to real images with higher S/Ns.
@@ 4 Results
The algorithm simultaneously processed the noisy simulations from the three Rubin bands and the Euclid VIS band, along with their respective PSFs. These noisy images served as initializations or first guesses for the algorithm. Subsequently, the algorithm iteratively minimized the loss functions 6-8, as detailed in Sect. 2.4. The algorithm was run for 200 iterations, with convergence typically observed within $50$-$100$ iterations for all images in our dataset. Figure 3 shows the convergence plot of the loss function for the deconvolved output in Fig. LABEL:subfig:mcdec1.
Figure 3: Loss function for the galaxy shown in Fig. LABEL:subfig:mcdec1. Convergence is guaranteed at around 100 iterations when the relative change in loss value is $<10^{-3}$ and the curve is flat.
@@ 4.1 Flux leakage test
As a validation to ensure no flux leakage between channels during the joint deconvolution of Rubin and Euclid images, we conducted a unit test. We assumed three distinct Gaussians placed separately in the Rubin channels, with the Euclid simulation being a weighted sum of these three images, resulting in three disjoint Gaussians. Post-deconvolution, the Gaussians remained intact without any structure extending beyond their boundaries. This confirms that the structure present in the deconvolved image within each Rubin band is independent of the structures in other bands and each image accurately retains only the information relevant to its specific band. Figure 4 provides a visual demonstration of these findings.
@@ 4.2 Deconvolved outputs
We present two examples of deconvolved images in Fig. LABEL:fig:deconv_outputs, illustrating the algorithm’s capability to recover features that were lost in the original Rubin simulations. Visually, the deconvolved outputs seem to capture the variations between different bands. Qualitatively, these outputs exhibit high quality with minimal background noise and result in clean residuals. However, in Sect. 4.3, we demonstrate that employing a deep-learning-based denoiser further enhances the already impressive results achieved initially. Table 2 presents the NMSE computed with respect to the ground-truth HST image for the pre-denoised and post-denoised images.
4.3 Deep-learning-based denoising
@@
After obtaining the deconvolved outputs, we feed them to DRUNet, a neural network proposed by Zhang et al. (2022) that combines U-Net (Ronneberger et al., 2015) and ResNet (He et al., 2016). U-Net is renowned for its efficiency in image-to-image translation, while ResNet excels in increasing modeling capacity through stacked residual blocks. Inspired by FFDNet (Zhang et al., 2018a), which incorporates a noise level map as input, DRUNet enhances U-Net by integrating residual blocks to improve prior denoising modeling. Similar approaches that combine U-Net and ResNet can be found in other studies (Zhang et al., 2018b; VenkateshG et al., 2018). The backbone of DRUNet is a U-Net architecture with four scales, and the schematic proposed by Zhang et al. (2022) is depicted in Fig. 5.
@@
We chose DRUNet because it is a non-blind denoiser, meaning it takes the noise map as input. This ensures that any unknown noise can be estimated and given as input for denoising. We estimate the noise level map in our deconvolved images using the scikit-image package (van der Walt et al., 2014) by calculating the average noise level within four $16\times 16$ pixel squares placed at each corner of the image. This approach ensures that only background noise is measured, avoiding any contribution from the signal. This map is then fed to the pre-trained DRUNet, along with the deconvolved image. The denoised outputs are illustrated in Fig. LABEL:fig:denoise_outputs. It is observed that the denoiser exclusively eliminates noise from the background without affecting the main structure of the galaxies. Although the enhancement in image quality is marginal (since the original deconvolved image is already of high quality), the NMSE computed with respect to the ground-truth HST image decreases, as shown in Table 2. The most notable improvement is observed in the $z$-band images. Even though DRUNet was trained on a combination of images from the BSD (Chen & Pock, 2017), Waterloo Exploration Database (Ma et al., 2017), DIV2K (Agustsson & Timofte, 2017), and Flick2K (Lim et al., 2017), it is remarkable that it works well on astronomical data, showing great generalization. Finally, the fractional error in the output flux as a function of the $i$-band magnitude is shown in Fig. 6, which indicates that the mean flux error is less than $5\%$ for the entire magnitude range for all the bands.
@@
Figure 5: DRUNet architecture, which incorporates an additional noise level map as input and integrates U-Net (Ronneberger et al., 2015) with ResNet (He et al., 2016). ”SConv” stands for strided convolution, and ”TConv” stands for transposed convolution. Image credits: Zhang et al. (2022).
@@
Figure 6: Fractional error in the output flux as a function of the $i$-band magnitude (which is chosen in order to have the same scale on the x-axes). The dots correspond to the individual galaxies, and the gray line is the best-fit line after binning the magnitude values.
@@
NMSE $r$ band $i$ band $z$ band
@@
 [^2] The NMSE is calculated for the pre-denoised and post-denoised images with respect to the ground-truth HST image.
4.4 Improvement compared to an independent deconvolution of each band
To demonstrate the advantages of jointly deconvolving multiple photometric bands, we performed independent deconvolution of each band using the deconvolution framework introduced in Akhaury et al. (2024). The method proceeds in two steps: it first de-convolves the input using Tikhonov regularization (Tikhonov & Arsenin, 1977), and denoises the result using SUNet (Fan et al., 2022), a state-of-the-art Swin transformer-based architecture. The results, presented in Fig. 8, show that the joint deconvolution method outperforms the independent deconvolution of individual bands. The joint method enables us to leverage the correlation between the different bands and the space-based image, thus improving the final output.
@@ 5 Conclusion
We have presented a novel image deconvolution technique designed to improve the resolution of multiband ground-based data by leveraging higher-resolution space-based observations. Our approach, which focuses on the joint deconvolution of Rubin and Euclid images, effectively exploits the overlapping spectral coverage of the Rubin $r$, $i$, and $z$ bands with the Euclid VIS band. Through rigorous testing, we have demonstrated that our iterative algorithm successfully recovers fine details while preserving the flux for each band. Different noise levels were tested, and the resolution achieved by ground-based data is close to that of HST. Our results indicate that a joint deconvolution of all data outperforms independent deconvolutions of individual photometric bands using existing state-of-the-art methods. By utilizing the Euclid VIS-band image as a term that provides additional information, our technique ensures that the deconvolved Rubin images retain high spatial resolution and accurate photometric measurements. The integration of deep-learning-based denoising using DRUNet enhances the quality of the deconvolved output, further reducing background noise without altering the main structures of the galaxies.
The potential applications of our method extend beyond the Euclid-Rubin pair, offering a versatile solution to improving the resolution of ground-based images in multiple photometric bands as long as there exists a space-based image of the same field of view in a band that encompasses all ground-based filters. In the future, we intend to test our deconvolution method on images of the Perseus cluster by using ground-based observations from the Canada–France–Hawaii Telescope (CFHT) and space-based observations from the Euclid Early Release Observations (ERO) public release.
Data availability
@@ Acknowledgements.
Figure 8: Galaxy shown in Fig. LABEL:subfig:mcdec2 deconvolved using two different approaches, illustrating that joint deconvolution outperforms independent deconvolutions of individual photometric bands. The joint method allows us to leverage the correlation between the different bands and the space-based image, thus improving the final output. First column: Euclid VIS image. Second: Rubin simulations in the $r$, $i$, and $z$ bands. Third: Independently deconvolved SUNet outputs for the three bands. Fourth: Corresponding joint deconvolution outputs followed by denoising with DRUNet. Fifth: Ground-truth HST images.
Appendix A Generalization to objects with non-flat SEDs
To assess the impact of information transfer across different bands for objects with a non-flat SED, we conducted an experiment where we successively replaced the Rubin bands with pure noise, ensuring that no galaxy signal was present. We then analyzed how features from the high-resolution Euclid image propagated into the deconvolved Rubin bands. The outputs are shown in Fig. LABEL:fig:noise_test.
Our results confirm that reconstructed features appear only in bands where the original galaxy signal is present. This demonstrates that the algorithm does not artificially imprint Euclid information onto Rubin bands lacking real data. In cases where a structure is visible in the Euclid image but absent from one or more Rubin bands, the problem becomes degenerate: the feature could either be entirely attributed to a single band or distributed across multiple bands. However, our findings indicate that the outputs are directly influenced by the input data in each band rather than being dictated solely by the high-resolution Euclid image. Our results demonstrate that the joint deconvolution method effectively utilizes the available signal in each band while respecting the constraints imposed by the data. This also confirms that the method would work for objects with non-flat SEDs, where signal may be present in only one band but absent in others, ensuring that features are accurately transferred according to their actual distribution across the bands.
As mentioned in Sect. 4, we find that the algorithm converges within $200$ iterations when the signal is available in all three bands. When one band is replaced with a noise map, as shown in Fig. LABEL:subfig:noise_test1, convergence requires approximately $1000$ iterations. Replacing a second band with a noise map, as seen in Fig. LABEL:subfig:noise_test2, further increases the iteration count to around $5000$. This demonstrates that incorporating more data across different bands significantly accelerates the convergence of the loss functions as the algorithm can benefit by capturing the correlations between these bands.